# Nonlocal interaction and quantum friction in sliding Bi2Se3 topological surfaces

### v3

@@ Acknowledgement
We gratefully acknowledge Qiang Sun for valuable discussions. This research was supported by Yunnan Fundamental Research Projects (Grant No. 202401AU070125), the Special Basic Cooperative Research Programs of Yunnan Provincial Undergraduate Universities’ Association (Grant No. 202301BA070001-114), Yunnan Provincial Department of Education Science Research Fund Project(No. 2025J0942), 2025 Self-funded Science and Technology Projects of Chuxiong Prefecture(Grant No. cxzc2025004, cxzc2025008), Chuxiong Normal University Doctoral Research Initiation Fund Project(No. BSQD2407), Yunnan Provincial Department of Education Science Research Fund Project(No. 2025J0942) and Dongying Science Development Fund Project(No. DJB2023015).
(a) The imaginary part of quantum action as a function of $v$, for the case $v_{\mathrm{F}}=3\times 10^{5}$$4.2\times 10^{5}$ and $2.2\times 10^{5}$, in units of $g^{2}TV$, with (a) $m=0.01$, (b) $m=1$, (c) $m=10$ and (d) $m=100$.
(b) The dissipative force as a function of $v$, for the case $v_{\mathrm{F}}=3\times 10^{5}$$4.2\times 10^{5}$ and $2.2\times 10^{5}$, in units of $g^{2}TV$, with (a) $m=0.01$, (b) $m=1$, (c) $m=10$ and (d) $m=100$.

### v5

@@ Acknowledgement
We gratefully acknowledge Qiang Sun for valuable discussions. This research was supported by Yunnan Fundamental Research Projects (Grant No. 202401AU070125), the Special Basic Cooperative Research Programs of Yunnan Provincial Undergraduate Universities’ Association (Grant No. 202301BA070001-114), Yunnan Provincial Department of Education Science Research Fund Project (No. 2025J0942), 2025 Self-funded Science and Technology Projects of Chuxiong Prefecture (Grant No. cxzc2025004, cxzc2025008), Chuxiong Normal University Doctoral Research Initiation Fund Project (No. BSQD2407, BSQD2507), Yunnan Provincial Department of Education Science Research Fund Project (No. 2025J0942) and Dongying Science Development Fund Project (No. DJB2023015).